If a body covers first half of its journey with a uniform spees v1 and the second half of journey in a uniform speed v2 ;then the average speed of the body is 
 
a]v1+v2                                    b]2v1 multiplied by v2 divided by                     
                                                  v1+v2 
 
c]v1v2 divided by v1+v2           d]v1 multiplied by v2
 
The correct answer is the second option but i diddnt undsrstand hoe the 2 into v1 v2 came there

Grade:11

1 Answers

Arun
4 years ago

Avg speed=

total distance covered/ total time taken

Suppose the total distance=s

Time taken for first half t1=distance/speed=(s/2)/v1

=S/2v1

Time taken for second half=s/2v2

Total time=s/2v1)+(s/2v2)=s/2(1/v1+1/v2)

=S(v1+v2)/2v1v2

Avg speed= S/ total time taken

By solving this you get

Avg speed=2v1v2/(v1+v2)